# Date of publication: December, 2025

### v1

@@ Abstract
microgrid, renewable energy, photovoltaic system, energy storage system, lithium-ion battery, lithium iron phosphate battery, high mountain, engineering.
@@ I Introduction
Energy supply in high mountains remains an open issue to be solved. Generally, grid connection is impractical, because the establishment and maintenance costs are too high. It also causes many environmental problems, e.g., the destruction of the natural ecology in the mountain. Diesel generators as energy suppliers have been used in high mountains for a long time despite their air pollution and noise problems. Due to the development of renewable energy (such as solar, wind, and hydropower), the usage of diesel generators is reduced, lowering the emissions of greenhouse gases (GHGs). However, owing to their fluctuating nature, most renewable energy sources exhibit intermittent features. To deal with the problem of renewable energy intermittency, energy storage systems (ESSs) are necessary [2, 3]. In the past, lead–acid batteries were heavily used as ESSs, accompanied by many environmental issues, e.g., poisoning, leaks, contamination of the environment, and damage to the ecosystem [4, 5]. Recently, lithium-ion (Li-ion) batteries [6, 7], as a greener alternative, have started to replace lead–acid batteries in ESSs.
Tianchi Lodge, a famous “five-star” mountain hut in Taiwan [8, 9], has adopted the off-grid solar energy storage system to provide electricity to the lodge since 2013. However, lead–acid batteries were implemented as ESS, accompanied by many environmental issues. In 2020, the lead–acid batteries were replaced by eco-friendly lithium iron phosphate (LFP) batteries. A better and greener off-grid solar energy storage system has been established.
@@ II-A Current status
Tianchi Lodge (coordinate: $24^{\circ}02^{\prime}43.15^{\prime\prime}$N, $121^{\circ}16^{\prime}46.77^{\prime\prime}$E) is an accommodation mountain lodge in Taiwan, located at an altitude of 2,860 m (9,380 ft) [10], at kilometer 13.1 of the Nenggao Cross-Ridge Historic Trail (as shown in Fig. 3). It is situated in a sheltered mountain valley on the west side of the North Peak of Nenggao, under the jurisdiction of the Nantou Branch, Forestry and Nature Conservation Agency, Ministry of Agriculture, Taiwan. Its administrative area is located in Ren’ai Township, Nantou County.
The current Tianchi Lodge utilizes natural lighting, has solar panels installed on the roof, generates electricity using renewable energy, and is equipped with a lithium-ion battery ESS [11]. Its electricity can provide hot water and nighttime lighting for the lodge. The specially designed sloping roof allows rain and snow to flow down smoothly. The lodge is near two waterfalls, so there is no shortage of water. It will not be short of water during the dry season. It is a truly environmentally friendly green building. The lodge also features a specially designed viewing platform on the second floor with large, operable glass windows, allowing visitors to enjoy the surrounding mountain scenery. The entrance corridor is spacious and includes a public social space with tea, tables, and chairs, providing hikers with a more comfortable activity space [12]. Due to these architectural designs and infrastructure, Tianchi Lodge has been dubbed a “five-star” luxury lodge in the mountain community by hikers [8, 9].
@@ II-B Brief history
The Forestry Bureau commissioned Wusiasyong Architects in 2003 and 2005 to plan and design the renovation of Tianchi Lodge [19, 20]. Tianchi Lodge was demolished and rebuilt on its original site in 2011. The project was constructed by Sinyuan Construction Co., Ltd. [21, 22] and supervised by Huangrueiming Architects [23, 24]. The concept of green building design was introduced during the reconstruction. It is a two-story steel structure building with a roof facing due south. It is equipped with solar water heating panels (solar water heating system), a solar power generation system, and uses natural materials such as wood and gravel to replace premixed concrete. It is constructed with an earthquake-resistant steel structure and fireproof paint [24]. It has rooms, a restaurant, a kitchen, a toilet, and a solar power generation system [25]. It is planned to have five eight-person rooms, twelve three-person rooms, and four six-person rooms, which can accommodate one hundred mountain enthusiasts. It was completed in June 2012  [25]. It opened to the public in 2013 [12]. The Tianchi Lodge construction project won the “2011-2012 Excellent Agricultural Construction Project Award (Architecture Category)” [24].
@@ III Architecture of energy system
Figure 2: Energy Architecture of Tianchi Lodge (2020). Energy input is provided in part by solar arrays and a diesel generator. The hybrid solar inverters directly supply power to the lodge’s 220 VAC load. The power of the 110 VAC load is obtained through a transformer (220 VAC / 110 VAC). An Energy storage system stores and releases intermittent renewable energy from solar arrays. A diesel generator is used as the backup power source.
@@ III-A Photovoltaic systems (PV systems)
A photovoltaic (PV) system converts sunlight directly into direct current (DC) electricity using solar panels.
In Tianchi Lodge, there are three PV arrays, providing a total power generation capacity of 11.2 kWp. The first channel of the PV array is rooftop-mounted on the second floor of the lodge with a power generation capacity of 7.84 kWp. The second channel of the PV array is rooftop-mounted on the first floor of the lodge with a power generation capacity of 1.68 kWp. The third channel of the PV array is rooftop-mounted on the toilet with a power generation capacity of 1.68 kWp.
@@ III-B Hybrid solar inverters
The hybrid solar inverter serves the crucial role of linking PV systems, ESS, AC loads, and backup power sources. The total nominal power of the hybrid solar inverter system is 15 kW.
@@ III-C LFP energy storage system (ESS)
ESS serves a role to balance the intermittent renewable energy.
@@ III-D Battery management system (BMS)
To use the battery safely, a well-designed BMS is required. A BMS is any electronic system that manages a battery cell, pack, or module system, such as protecting the battery from operating outside its safe operating area, monitoring its status, calculating secondary data, reporting that data, controlling its environment, and cell balancing it [36].
III-E Load
Two AC loads with different frequencies (220 VAC and 110 VAC) are used in the lodge.
@@ III-F Backup power system
In Tianchi Lodge, a three-phase diesel generator of nominal output voltage 220 VAC is installed on site. The output power capacity is 20 kW.
@@ IV Historical development of energy systems
In the past, lead-acid battery energy storage systems combined with diesel generators were commonly used for the power supply of mountain huts. However, the environmental damage caused by lead pollution from lead-acid batteries and their harm to human health have created serious problems for their use in high-altitude areas. Furthermore, lead-acid batteries will age and deteriorate after a few years of use, eventually becoming unable to store electricity.
@@
In terms of solar arrays, 1.68 kWp solar panels were installed on the roof of the first floor of the lodge, 7.84 kWp solar panels were installed on the roof of the second floor of the lodge, and 1.68 kWp solar panels were installed on the roof of the toilet. In terms of energy storage, the lodge had 86 kWh of lead-acid batteries, and the toilet had 9.6 kWh of lead-acid batteries. The inverter had an AC output capacity of 5 kW and a rated voltage of 220 VAC. The diesel generator had a power generation capacity of 20 kW and a rated voltage of 220 VAC.
@@ V Summary and Outlook
Using renewable energy and eco-friendly ESS in high mountains is not easy. It took eight years to reach the goal in Tianchi Lodge. Since 2013, renewable energy (PV systems) has been used. However, the lead-acid battery ESS still causes pollution in high mountains. Until 2020, the eco-friendly LFP ESS replaces the retired lead-acid battery ESS, a better and green off-grid solar energy storage system is build.
In the next step, we should install an EMS to check the energy input and output. Figuring out the energy usage of the lodge. Further suggestions can be given after the energy usage analysis, such as “Is it necessary to expand the PV system?” or “Is it necessary to expand the ESS?”

### v2

@@ Abstract
microgrid, renewable energy, photovoltaic system, energy storage system, lithium-ion battery, lithium iron phosphate battery, high mountain, mountain hut, high altitude, engineering.
@@ I Introduction
Energy supply in high mountains remains an open issue to be solved. Generally, grid connection is impractical, because the establishment and maintenance costs are too high. It also causes many environmental problems, e.g., the destruction of the natural ecology in the mountain. Diesel generators as energy suppliers have been used in high mountains for a long time despite their air pollution and noise problems. Due to the development of renewable energy (such as solar, wind, and hydropower), the usage of diesel generators is reduced, lowering the emissions of greenhouse gases (GHGs). However, owing to their fluctuating nature, most renewable energy sources exhibit intermittent features. To deal with the problem of renewable energy intermittency, energy storage systems (ESSs) are necessary [2, 3]. In the past, lead-acid batteries were heavily used as ESSs, accompanied by many environmental issues, e.g., poisoning, leaks, contamination of the environment, and damage to the ecosystem [4, 5]. Recently, lithium-ion (Li-ion) batteries [6, 7], as a greener alternative, have started to replace lead-acid batteries in ESSs.
Tianchi Lodge, a famous “five-star” mountain hut in Taiwan [8, 9], has adopted the off-grid solar energy storage system to provide electricity to the lodge since 2013. However, lead-acid batteries were implemented as ESS, accompanied by many environmental issues. In 2020, the lead-acid batteries were replaced by eco-friendly lithium iron phosphate (LFP) batteries. A better and greener off-grid solar energy storage system has been established.
@@ II-A Current status
Tianchi Lodge (coordinate: $24^{\circ}02^{\prime}43.15^{\prime\prime}$N, $121^{\circ}16^{\prime}46.77^{\prime\prime}$E) is an accommodation mountain lodge in Taiwan, located at an altitude of 2,860 m (9,380 ft) [10], at 13.1 km of the Nenggao Cross-Ridge Historic Trail (as shown in Fig. 1). It is situated in a sheltered mountain valley on the west side of the North Peak of Nenggao, under the jurisdiction of the Nantou Branch, Forestry and Nature Conservation Agency, Ministry of Agriculture, Taiwan. Its administrative area is located in Ren’ai Township, Nantou County.
The current Tianchi Lodge utilizes natural lighting, has solar panels installed on the roof, generates electricity using renewable energy, and is equipped with a lithium-ion battery ESS [11]. Its electricity can provide hot water and nighttime lighting for the lodge. The specially designed sloping roof allows rain and snow to flow down smoothly. The lodge is near two waterfalls, so there is no shortage of water. The water is sufficient even during the dry season. It is an environmentally friendly green building. The lodge also features a specially designed viewing platform on the second floor with large, operable glass windows, allowing visitors to enjoy the surrounding mountain scenery. The entrance corridor is spacious and includes a public social space with tea, tables, and chairs, providing hikers with a more comfortable activity space [12]. Due to these architectural designs and infrastructure, Tianchi Lodge has been dubbed a “five-star” luxury lodge in the mountain community by hikers [8, 9].
@@ II-B Brief history
The Forestry Bureau commissioned “Wu, Hsia-Hsiung Architect & Associates” in 2003 and 2005 to plan and design the renovation of Tianchi Lodge [19, 20]. Tianchi Lodge was demolished and rebuilt on its original site in 2011. The project was constructed by Sinyuan Construction Co., Ltd. [21, 22] and supervised by Huangrueiming Architects [23, 24]. The concept of green building design was introduced during the reconstruction. It is a two-story steel structure building with a roof facing due south. It is equipped with solar water heating panels (solar water heating system), a solar power generation system, and uses natural materials such as wood and gravel to replace premixed concrete. It is constructed with an earthquake-resistant steel structure and fireproof paint [24]. It has rooms, a restaurant, a kitchen, a toilet, and a solar power generation system [25]. It is planned to have five eight-person rooms, twelve three-person rooms, and four six-person rooms, which can accommodate one hundred mountain enthusiasts. It was completed in June 2012 [25] and opened to the public in 2013 [12]. The Tianchi Lodge construction project won the “2011-2012 Excellent Agricultural Construction Project Award (Architecture Category)” [24].
@@ III Architecture of energy system
Figure 2: Energy architecture of Tianchi Lodge (2020). Energy input is provided in part by PV arrays and a diesel generator. The hybrid solar inverters directly supply power to the lodge’s 220 VAC load. The power of the 110 VAC load is obtained through a transformer (220 VAC / 110 VAC). An ESS stores and releases intermittent renewable energy from PV arrays. A diesel generator is used as the backup power source.
@@ III-A Photovoltaic systems (PV systems)
A photovoltaic (PV) system converts sunlight directly into direct current (DC) electricity using solar panels. Then the electricity is sent to a hybrid solar inverter for further power conversion.
In Tianchi Lodge, there are three PV arrays, providing a total power generation capacity of 11.2 kWp. The first channel of the PV array is rooftop-mounted on the second floor of the lodge with a power generation capacity of 7.84 kWp. The second channel of the PV array is rooftop-mounted on the first floor of the lodge with a power generation capacity of 1.68 kWp. The third channel of the PV array is rooftop-mounted on the toilet with a power generation capacity of 1.68 kWp. In order to integrate the PS arrays into the newly installed hybrid solar inverter system, the series and parallel connections of the PV arrays must be readjusted to meet the inverter input specifications.
@@ III-B Hybrid solar inverters
The hybrid solar inverter serves the crucial role of linking PV systems, ESS, AC loads, and backup power sources. The total nominal power of the hybrid solar inverter system is 15 kW. Considering the challenges of maintaining mountain huts, a system design employing multiple units operating in parallel to simultaneously output power was chosen for reliability and stability. Under normal operating conditions, all units can supply power to the lodge. In the event of any failure, the faulty unit can be disconnected, and the remaining working units can continue supplying power while the system provider is notified for maintenance. During this process, the inverter units can still supply power to the lodge, significantly reducing downtime caused by system failures.
@@ III-C LFP energy storage system (ESS)
ESS is a technology that stores excess electricity and releases it when needed, acting as a “reservoir” in the power system. It is mainly used to stabilize the power grid, integrate intermittent renewable energy sources (such as solar, wind, and hydropower), and provide peak shaving and valley filling to improve power supply efficiency and reliability. Its core components include battery modules and BMS, and it regulates power supply and demand through charging and discharging.
ESSs in high-altitude areas have long used lead-acid batteries. However, the heavy metal lead and highly acidic electrolyte contained in these batteries can severely pollute soil and water sources and harm human health (nervous system, kidneys, and brain) if discarded or improperly disposed of. Without adequate equipment, the recycling and processing process can also generate pollutants such as lead dust and acid mist.
To enable the use of green energy in high-altitude areas, we need the support of new technologies. The vigorous development of electric vehicles (EVs) in recent years has also driven the progress of Li-ion battery technology. Compared with lead-acid batteries, which have a history of over a century but are highly polluting, Li-ion batteries have many advantages, such as high energy density, light weight, no memory effect, low self-discharge rate, and long lifespan. Li-ion batteries have been widely used in 3C products and EVs. The 2019 Nobel Prize in Chemistry was awarded to John B. Goodenough, M. Stanley Whittingham, and Akira Yoshino for their contributions to the development of Li-ion batteries  [37, 38, 39].
Through long-term technological improvements, lithium-ion batteries have evolved into many different types (such as ternary Li-ion batteries and LFP batteries), each with its own advantages and disadvantages. For example, ternary Li-ion batteries have high energy density but slightly weaker safety, while LFP batteries, although having slightly lower energy density, are safe, stable, and have a long cycle life. To address the challenges of maintenance in high-altitude environments, safe and highly stable LFP batteries are chosen as the energy storage system.
@@ III-D Battery management system (BMS)
To use the battery safely, a well-designed battery management system (BMS) is required. A BMS is any electronic system that manages a battery cell, pack, or module system, such as ensuring the battery operating within its safe operating area, monitoring its status, calculating secondary data, reporting that data, controlling its environment, and cell balancing [36].
III-E Loads
Two AC load channels with different voltages (220 VAC and 110 VAC) and the same frequency of 60 Hz are used in the lodge. When designing the system, it is necessary to assess the power consumption of the lodge during normal operation and during high-load operation.
@@ III-F Backup power system
In Tianchi Lodge, a three-phase diesel generator with nominal output voltage 220 VAC and frequency of 60 Hz is installed on site. The output power capacity is 20 kW.
@@ IV Historical development of energy systems
In the past, lead-acid battery ESSs combined with diesel generators were commonly used for the power supply of mountain huts. However, the environmental damage caused by lead pollution from lead-acid batteries and their harm to human health have created serious problems for their use in high-altitude areas. Furthermore, lead-acid batteries will age and deteriorate after a few years of use, eventually becoming unable to store electricity.
@@
In terms of PV arrays, 1.68 kWp solar panels were installed on the roof of the first floor of the lodge, 7.84 kWp solar panels were installed on the roof of the second floor of the lodge, and 1.68 kWp solar panels were installed on the roof of the toilet. In terms of energy storage, the lodge had 86 kWh of lead-acid batteries, and the toilet had 9.6 kWh of lead-acid batteries. The inverter had an AC output capacity of 5 kW and a rated voltage of 220 VAC. The diesel generator had a power generation capacity of 20 kW and a rated voltage of 220 VAC.
@@ V Summary and Outlook
Using renewable energy and eco-friendly ESS in high mountains is not easy. It took eight years to reach the goal in Tianchi Lodge. Since 2013, renewable energy (PV systems) has been used. However, the lead-acid battery ESS still causes pollution in high mountains. Until 2020, the eco-friendly LFP ESS replaces the retired lead-acid battery ESS, a better and green off-grid solar energy storage system is built. After years of use, its operational reliability has been verified.
In the next step, we should install an EMS to check the energy input and output. Figuring out the energy usage of the lodge. Further suggestions can be given after the energy usage analysis, such as “Is the solar power generation sufficient, and is it necessary to expand the solar power system?” or “How many days can the ESS withstand without solar power generation, and is it necessary to expand the ESS?” Although diesel generators serve as a backup solution for the backup energy system, we still hope that their usage time can be as short as possible to minimize their damage to the high-altitude environment.